as A Questions 1-4 refers to the five titration curves (labelled mL of titrant mL of titrant ml of titrant mL of titrant mL of titrant 1 Which curve shows the titration of a strong acid with a strong base? AA BB CC DD EE Fnone of the above 2 Which curve shows the titration of a weak acid with a strong base? A A BB CC DD EE Fnone of the above 3 Which curve shows the titration of a polyprotic base with a strong acid? AA BB CC DD EE Fnone of the above 4 Which curve shows the titration of a weak base with a strong acid? AA BB CC DD EE Fnone of the above 5 A certain weak acid has a pKa of 4.0. When the pH pKa, the [H] is closest to A 4X10M B 10 M C 4.0 M D 1x10 4 M E 104 M F¼M 6 If the pH of a solution is 3, the [H] is closest to: A 3M 80.3M C 0.003M D 1 X10M E 1 X 10 M F 1 X 10 14M G 3 X 103M 7 What is the pH of a 0.01M weak acid whose Ka 104HIA/[HA] A 1.0 B 2.0 C 3.0 D 4.0 E 5.0 F 7.0 G 11.0 H-4 8 If the (H] of a solution is 10 °M, the pH is and the solution is A 5; basic B 5;acidic C 9; acidic D 9; basic E 7;neutral F none of the
Answer #1

1•] For a strong acid- strong base titration equivalence point should be at pH=7. Hence answer can be A or B. But here strong acid is the analyte and strong base is the titrant .Hence the answer is B.

2•] weak acid vs strong base gives equivalence point in the basic region that is above pH 7. Answer is D.

3•]. Polyprotic base vs strong acid titration curve has more than one equivalence points. Hence answer is E.

4•] .

Weak base vs strong acid titration curve has equivalence point in acidic region ,pH<7. Hence C.

5•]. pH= -log[H+]

[H+]=10^(-pH)

=10^(-4)

=0.0001 M

Answer D.

6•]. [H+] =10^(-pH)

=10^(-3)

=0.001 M

Answer D.

7•]. For a weak acid

[H+] =?(ka÷c)

=?{(10^-4)÷0.01}

=0.1000 M

pH= -log[0.1000]

=1

answer A

8•] pH= -log[10^-5]

=5

Hence the solution is acidic. Answer B.